# Cultural Transmission, Technology, and Treatment of the Elderly

### v1

@@ 3.1.3 Knowledge and the Elderly
The elderly often provide valuable services securing them a livelihood, and often provide administrative services, which can convey power and prestige to the elderly. Among the less developed peoples in Simmons’ sample, the elderly may receive payment for the provision of services that confer something of material value, such as childcare or food processing, or even technical knowledge or administrative know-how. The status of the elderly increases and treatment improves with the importance of tradition in society. Elderly status also improves with the amount of specialized knowledge the elderly have.
But Simmons also discusses how the elderly also often provide purely cultural goods and services in exchange for compensation. For example, elderly Navajo were compensated for their knowledge of magic songs and charms, while among the Pomo and the Lengua, elderly medicine men earned fees for performing important rites (Simmons, 1970, p. 37-38). Thus, some types of property and associated property rights appear to be established by cultural conditioning, as is the case with rights over songs and charms, while at the same time, other types of cultural conditioning alter or expand property rights in ways that aid the elderly. As an example, rules for sharing game or food tabooing might be thought of in this way.
The belief that the elderly possess magical powers is common among subsistence hunters and hunters, instilling fear and awe of the elderly. The elderly are often believed to be endowed with the power to bring and cure disease, and these beliefs are actively cultivated. The cultures in which such beliefs were prevalent operate in widely varied material circumstances; the Yahgan (Tierra del Fuego), the Arunta (Australia), the Greenland Inuit, the Pomo (California Desert), and the Witoto (Northern Peru). Rosenberg (2009, p. 34) notes that the Ju/’hoansi hold the belief that some elderly have powers of healing, which are due to their experience allowing them to use certain mystical forces without being overwhelmed by them.
@@ 3.1.4 Inculcation
Attitudes about the elderly and cultural practices directed towards the elderly are consciously nurtured in the young, sometimes at considerable cost to parents. Simmons (1962) is clear to indicate that these forms of elderly treatment arise due to the application of costly resources and effort. Simmons notes that “Respect for old age has resulted from social discipline,” (Simmons, 1970, p. 50), and, writing more expansively (his italics):
“…while human beings can be as indifferent to the dependency needs of their ancestors…they can also learn, be taught, inspired, or impelled to respect, succor, and sustain their elders…these patterns of group and filial responsibility are ‘passed on’ from generation to generation…This learned behavior is called culture.” (Simmons, 1962, p. 37)
Thus, some degree of cultural conditioning allows the elderly to consume and control resources that they would not be able to otherwise.
@@ 3.2 Elderly Status and Cultural Complexity
Similar conclusions also emerge in subsequent cross-cultural research on the treatment of the elderly. Press and McKool Jr. (1972, p. 298) give a brief summary of the “ostensibly universal generalizations about status of the aged” offered by Cowgill and Holmes (1971), which, more or less, follow those of Simmons (1970). These generalizations include that the status of the elderly is high when knowledge accumulation is important in the economy, when the elderly can hold property rights, and when the elderly can perform economically meaningful tasks. In fact, Press and McKool Jr. (1972, p. 299), in developing their own list of determinants of elderly status, came to the conclusion that (their italics) “…our independently derived determinants coincided with Cowgill’s.”
Our discussion has so far focused on the elderly among hunter-gatherers and cultures that rely on subsistence agriculture or animal husbandry for subsistence, but much subsequent cross-societal research has focused on how elderly treatment evolves as societies progress further along some dimension of technological complexity, which seems to encompass dimensions such as the definition of property rights, social complexity, and technological change in the basic mode of production of the economy. In this research, elderly treatment becomes enmeshed in research on the complexity of familial structure, but a broad trend in elderly treatment that has emerged from this line of inquiry is that it is difficult to define an exact trajectory of elderly treatment as the economy becomes more complex. Put differently, elderly treatment varies in a non-linear fashion with the level of development of society. One difficulty plaguing this branch of research is the reliance on overly vague terminology, where terms such as “status” and “good treatment” are used without formal definition.131313 Dowd (1981) and Balkwell and Balswick (1981) note the problems caused by the poor definition of relevant terms. What does it mean to say, for example, that the elderly are treated well in a subsistence hunting and gathering community but not all that well treated in a modern industrialized economy? Some researchers may claim that the status of the elderly is increasing in the level of societal development, and while other researchers may disagree with this, the disagreement may really be about how status is defined and what the relevant range of development is.
Despite some confusion about these terms, an important characterization is due to the so-called “curvilinear hypothesis,”(Blumberg and Winch, 1972) as originally developed in Nimkoff and Middleton (1960).141414(Lee, 1996) is a good assessment and an overview of the hypothesis. See also Lee and Kezis (1981) and Baker and Miceli (2005). According to the hypothesis, there is a nonlinear relationship between social complexity and the extent of the family. Those societies at the lowest levels of development (that is, hunter-gatherer societies) and those at higher levels of development (intensive agriculturalists and industrial societies) tend to have smaller families. Those societies “in the middle” tend to have the most expansive family structures, and, as Lee (1979) documents, the elderly tend to have a high status in these familial structures. This literature considers the nexus of things documenting the importance of familial structure in elderly treatment, but often confounds what are really two different things: the elderly being treated preferentially because of a cultural predilection to do so, and the elderly commanding respect because they have accumulated resources in the past. This is at the crux of the matter: How do accumulation of assets and cultural transmission relate to each other?
@@ 3.3 Modernization and the Elderly
As globalization and technological change continue, as population growth slows, and as the global population ages, it is important to understand how technical change interacts with cultural practices relevant to the elderly. Although there seems to be some reason to believe that industrialization is detrimental to the well-being of the elderly (Dowd, 1981, e.g.), the relationship between different aspects of industrialization and the well-being of the elderly has turned out to be complex. Understanding these relationships is even important in modern societies, and the study of such problems naturally leads one to wonder how the institutions and cultural practices that provide care for the elderly are impacted by modern institutions, such as a social security system. These questions become more important when considering that much of the world’s population will be transitioning from lower to higher levels of development in the coming decades. Koda and Uruyos (2018) for example, find evidence that public transfers crowd out private care for the elderly and also that the development process, which leads to lower fertility and smaller families, tends to reduce family support for the elderly, offering some support for the theories presented in Koda and Uruyos (2018) and Cigno (1993). In contrast, Ko and Möring (2021) find in their study of pensions in rural China that receiving a public pension accompanies an increase in both well-being and the degree of intergenerational exchange. Some research (Kodate and Timonen, 2017) has described some of the complexity of the interaction between public funding, paid health care workers, culture, and family care.
Therefore, elderly treatment and status interact with the standard desire to smooth consumption in rather complex ways. In general, the desire to push consumption of current resources into the future may result in improving treatment through cultural means, accumulation of assets over time, or both. How exactly this happens depends on the nature of production and other basic characteristics of the economy.
Developing a model in which these trade-offs can be explored turns out to shed much light on these issues and maybe even lends insight into how elderly treatment will evolve in the coming decades. A model should allow us to better understand the interplay between elderly treatment and the underlying attributes of the economy and also generate predictions that capture some of the complexity in the ethnographic record. In the next section, we present such a model. An additional useful aspect of our model is that it requires defining some of the often used concepts in the empirical literature that remain vague, such as ‘status’ or ‘prestige’, or even what it means for the elderly to be ‘treated well.’
@@ 4 Theoretical Framework
The model is a standard infinite-horizon overlapping generations model, where at any given time young, middle-aged, and elderly agents co-exist. Agents make decisions solely in middle age. When young, agents do not earn any utility but can be inculcated to care about culture. In our base model, current middle-ageds cultivate a taste for giving gifts to the elderly in the young. When young people reach middle age, they are willing to transfer resources to the elderly, where their willingness to transfer resources and earn ‘warm glow utility’ as in Andreoni (1989) depends on the degree to which appreciation was cultivated in agents currently in middle age when they were young. [^15] In this respect, our model resembles one in which cultural heritage and society’s devotion to it is endogenous, in a similar fashion to Cozzi (1998). Cozzi’s model, however, is built on the premise that beliefs about the value of culture are self-fulfilling and are not actively cultivated.
The resulting model of endogenous transfers to the elderly is observationally equivalent to a model in which appreciation for a cultural good is cultivated in the young and the cultural good is supplied by the elderly. Thus, the endogenous transfer, warm-glow formulation of the model captures an element of the degree to which cultural transmission occurs and allows us to measure how important transfer of culture is in maintaining the elderly.
@@
The consumption of middle-aged and elderly people at time $t$ are $c_{mt}$ and $c_{et}$, while current gifts to the elderly are given by $g_{mt}$. Lifetime utility of a current middle ager depends upon consumption and gifts according to the log-linear utility function:
$U_{t}=(1-\eta_{t})\ln{c}_{mt}+\eta_{t}\ln g_{t}+\beta\ln c_{et+1}$ (1)
@@
In the lifetime utility specification (1), $\eta_{t}$ measures the relative importance of gift-giving in middle-aged utility.
Current middle-aged agents receive income $y_{mt}$ in the form of a numeraire. A middle-aged agent divides this income between consumption $c_{mt}$, savings $s_{mt}$, and gifts to the elderly $g_{mt}$. Income can also be used to teach children to appreciate culture. We model this ‘inculcative investment’ as a choice of the future value of the demand parameter $\eta_{t+1}$, that is, the value of $\eta$ that current children will have when they are middle-aged. [^16] As we noted previously, this distinguishes our model from the models of Olivera (2013) and Varvarigos (2021), which assume that inculcative investments increase the likelihood that children will have the same preferences as adults. One might refer to those models as ‘Simpsonian’ models reflecting Homer Simpson’s attitudes about parenting: “Kids are the best… You can teach them to hate the things you hate.” (The Simpsons, season 11 episode 233, 1999) The cost of setting this value in current resources is $d_{t}\eta_{t+1}$. We shall typically assume that $d_{t}=\delta y_{mt}$, where $\delta\in(0,1)$ so that inculcative costs are scaled to current income. This could be considered a simple way to capture the idea that the costs of inculcation are measured in opportunity costs, and when inculcation occurs, the value of the resources employed to inculcate is measured in units of income. [^17] Without this assumption equilibria would be non-stationary and analysis of balanced-growth paths would be impossible. Middle-aged consumption $c_{mt}$ is then income, less inculcative investment costs, savings, and gifts to the elderly, as follows.
$c_{mt}=y_{mt}-d_{t}\eta_{t+1}-s_{mt}-g_{mt}$ (2)
We now describe the components of $c_{et}$ to complete a description of lifetime utility (1). The elderly may also earn income, which we denote as $y_{et}$. $y_{et}$ might include the income earned by the elderly if they perform useful tasks or other functions, including administrative, advisory, or other types of service, as documented in Section 3. A broader interpretation of $y_{et}$ includes government policies such as social security or other government transfer payments in applications of the model to modern societies. In a hunter-gatherer society, the income $y_{et}$ received by the elderly could also derive from the productive efforts of others in the form of customary sharing of output.
Elderly consumption differs from the income of the elderly and can be increased by consumption derived from gifts received $g_{et}$. Elderly consumption may also derive from the usual source in an overlapping-generations model: returns from assets accumulated earlier in life. Elderly consumption at time $t$ then consists of returns from savings in middle age, $R_{t}s_{mt-1}$, income earned $y_{et}$, and gifts received:
@@
One point that requires clarification is how exactly this model of intergenerational gift giving relates more broadly to the transmission of culture. Shortly, we shall show that the model characterized by the utility function in (1) is observationally equivalent to a model in which cultural goods of endogenously determined value are transferred between generations.
@@ 4.1 Simple model mechanics
We begin by tackling a simplified version of the model in which there is no savings ($s_{mt}=0$ in Equation (3)). This specialization of the model serves three purposes. First, it familiarizes the reader with simple model intuition, isolating only the inculcation and gift decisions of current middle-aged agents. Second, it serves an empirical purpose in that this version of the model better characterizes premodern economies where asset formation and capital accumulation are not important, and most output depends on labor along with natural resources such as land to produce output, as is the case among hunter-gatherer societies and subsistence agricultural societies. Third, this model specialization is well suited to develop some insight into how property rights, inculcation, and elderly treatment are interrelated.
Substituting middle-aged consumption (2) and elderly consumption (3) into (1), and setting $s_{mt}=0$ gives:
@@
To simplify things even further, we assume for the time being that $\eta_{t}\in\{0,\eta\},0<\eta<1$ so that the current middle-aged generation makes a simple binary decision about whether to inculcate children to enjoy giving gifts to the elderly at fixed intensity $\eta$ at cost $d_{t}$ or not, while at the same time choosing a gift to give to the current elderly based on the previous generation’s inculcation decision. These assumptions allow us to solve the model as one would a stationary, infinite-horizon game in which each generation sequentially decides whether or not to inculcate the youth, (correctly) anticipating that future generations will maintain the (subgame-perfect Nash equilibrium) strategy.
Middle-aged lifetime utility when inculcation has occurred in the past, is continued in the present, and is expected by current agents to continue in the future is:
@@
Where $g_{et+1}$ is the gift expected in the next period, given that current youth are inculcated. We can see if continuing this inculcative equilibrium is stable by contemplating what happens when a current middle-aged agent deviates and does not engage in inculcative investment. A deviation from the equilibrium path, thereby discontinuing the inculcation policy, would result in no future gifts, so the current middle-aged agent would attain lifetime utility:
$U_{mt,0}=(1-\eta)\ln(y_{mt}-g^{\prime}_{mt})+\eta\ln g^{\prime}_{mt}+\beta\ln(y_{et+1})$ (6)
In the case where gift giving and inculcation is continued, we find optimal gifts by differentiating (5) with respect to $g_{mt}$ and solving gives:
@@
The current middle-aged agent therefore gives a gift proportion to disposable income, which is just income left over after inculcation. Having solved for equilibrium gifts given, we can easily solve for gifts received, $g^{*}_{et}$. As the relative size of the elderly population plays an important part in many discussions of the role of the elderly in the economy, we wish to include population changes in the model. We do this simply by assuming that the population grows (or declines) at a constant rate $n$, so $n_{t+1}=(1+n)n_{t}$. This implies that for gifts given to be equal in quantity to gifts received, so in equilibrium, $g^{*}_{et}=(1+n)g^{*}_{mt}$.
The middle-aged agent who discontinues the inculcation policy gives what amounts to a final gift of $g^{\prime}_{mt}=\eta y_{mt}$, and then receives no gift when elderly, in the following period.

### v2

@@ 3.1.3 Knowledge and the Elderly
The elderly often secure livelihoods by providing valuable services, including administrative tasks that convey power and prestige. Among less developed societies in Simmons’ sample, elders may receive payment for childcare, food processing, technical knowledge, or administrative expertise. Their status improves with the importance of tradition and the degree of specialized knowledge they possess.
Simmons also highlights the provision of cultural goods by the elderly in exchange for compensation. For instance, elderly Navajo were paid for knowledge of magic songs and charms, while Pomo and Lengua medicine men earned fees for performing rites (Simmons, 1970, p. 37-38). Cultural conditioning plays a role in establishing property rights over such intangible assets, as with rights over songs and charms, and in shaping practices like food tabooing or game-sharing rules that benefit the elderly.
Beliefs in the magical powers of the elderly are widespread among subsistence hunters and gatherers, instilling fear and respect. Elders are often thought capable of bringing or curing disease, beliefs actively cultivated across varied contexts, including the Yahgan (Tierra del Fuego), Arunta (Australia), Greenland Inuit, Pomo (California Desert), and Witoto (Northern Peru). Among the Ju/’hoansi, some elders are believed to have healing powers derived from their ability to harness mystical forces without being overwhelmed (Rosenberg, 2009, p. 34).
@@ 3.1.4 Inculcation
Attitudes and practices toward the elderly are consciously instilled in the young, often at significant cost to parents. Simmons (1962) emphasizes that such treatment arises from deliberate application of resources and effort, stating, ‘Respect for old age has resulted from social discipline’ (Simmons, 1970, p. 50). Expanding on this point, Simmons observes:
…while human beings can be as indifferent to the dependency needs of their ancestors…they *can also* learn, be taught, inspired, or impelled to respect, succor, and sustain their elders…these patterns of group and filial responsibility are ‘passed on’ from generation to generation…This learned behavior is called ‘culture’.” (Simmons, 1962, p. 37)
Cultural conditioning thus enables the elderly to access and control resources they might otherwise lack.
@@ 3.2 Elderly Status and Cultural Complexity
Subsequent cross-cultural research on elderly treatment reaches similar conclusions. Press and McKool Jr. (1972, p. 298) summarize Cowgill and Holmes (1971) and their “ostensibly universal generalizations about the status of the aged,” that align closely with Simmons (1970). These include the high status of the elderly in contexts where knowledge accumulation is economically important, property rights are accessible, and the elderly can perform meaningful economic tasks. In fact, Press and McKool Jr. (1972, p. 299) note that their independently derived determinants of elderly status coincided with Cowgill’s findings.
Although much focus has been on hunter-gatherers and subsistence agricultural or pastoral societies, later research explores how elderly treatment evolves with increasing technological complexity. This involves changes in property rights, social complexity, and production methods. However, defining a consistent trajectory for elderly treatment as societies develop is challenging. Research suggests a nonlinear relationship, with treatment varying across levels of societal complexity. Terms like ‘status’ and ‘good treatment’ are often poorly defined, complicating analysis. [^5] Dowd (1981) and Balkwell and Balswick (1981) highlight the issues with vague terminology. For example, comparisons of elderly treatment in hunter-gatherer versus modern industrial societies may hinge on unclear definitions of status and modernity.
A significant characterization is the ‘curvilinear hypothesis’ (Blumberg and Winch, 1972), as developed by Nimkoff and Middleton (1960). [^6] See Lee (1996) for an overview, as well as Lee and Kezis (1981) and Baker and Miceli (2005). This hypothesis posits a nonlinear relationship between social complexity and family structure: hunter-gatherers and advanced societies tend to have smaller families, while intermediate-level societies feature expansive familial structures. In societies of intermediate complexity, the elderly often hold high status, as documented by Lee (1979). However, this literature often conflates two factors: preferential treatment due to cultural norms and respect derived from accumulated resources. The core question remains: how do asset accumulation and cultural transmission interact?
3.3 Some quantitative evidence
Simmons (1970) provides extensive cross-tabulations of variables to highlight empirical regularities in elderly treatment. This section summarizes his data and presents evidence on the roles, inculcation, and treatment of the elderly across societies. Simmons’ data, while valuable, poses challenges for advanced analysis due to its small sample size and missing values. To address this, we constructed indices summarizing key societal features related to the elderly. These indices suggest meaningful relationships among societal characteristics. Appendix A provides a detailed discussion of the data and index construction.
The data includes qualitative traits scaled into indices, summarized in Table 1. The first four indices consider the elderly’s role in directly economically productive activities, in social production, in advice and knowledge provision, and the holding of titular roles in the society. The fifth index combines the first three forms of production. The sixth index considers the openness of the society to market-based economic activity. The seventh index measures positive treatment of the elderly, and the eighth index measures the amount of positive inculcation towards the elderly.
Index Description
Contribution to direct production Tasks such as trading, childcare, hunting support, agriculture, household work, and collecting fees for services.
Contribution to social production Performing marriages, initiation rites, funerals, and roles as priests or shamans.
Knowledge and advice provision Providing advice and information.
Titular duties Holding roles as chiefs, elders, and judges.
Combined production Sum of direct production, social production, and knowledge and advice provision indices.
Openness of economy Presence of property rights, trading, and currency usage.
Positive treatment Respect, distinction, and familial support for the elderly.
Positive inculcation Beliefs portraying the elderly as daimons, heroes, and/or children’s friends.
Table 1: Indices constructed from data in Simmons (1970).
We then correlated the six production and openness of economy indices with the two indices for positive treatment of the elderly and positive inculcation towards the elderly. Table 2 presents these correlations. Positive treatment of the elderly is positively correlated with the elderly’s contribution to production and with their provision of advice and knowledge, as well as with the combined production index and with their holding of titular duties. Positive inculcation regarding the elderly is also positively correlated with these categories except for titular duties. Positive treatment and positive inculcation are positively correlated. Interestingly, the open economy index is not correlated with positive treatment and shows some evidence (though not at standard statistical significance) or negative correlation with positive inculcation. We shall consider these patterns in the context of consistency with our formal model later in the paper.
Positive treatment Positive inculcation
Contribution to direct production $0.38^{*}$ $0.24^{*}$
Contribution to social production 0.21 $0.22$
Knowledge and advice provision $0.43^{*}$ $0.45^{*}$
Combined production $0.41^{*}$ $0.33^{*}$
Titular duties $0.36^{*}$ $0.15$
Openness of economy 0.05 -0.16
Positive treatment - $0.45^{*}$
Table 2: Pairwise correlations between indices based on Simmons (1970). denotes significance at 95%. See Appendix A for variable details.
@@ 3.4 Modernization and the Elderly
As globalization and technological change advance, population growth slows, and the global population ages, understanding the interaction between technical change and cultural practices related to the elderly becomes increasingly critical. While industrialization is often considered detrimental to elderly well-being (Dowd, 1981, e.g.,), the relationship between industrialization and elderly welfare is nuanced. This complexity remains relevant even in modern societies, raising questions about how institutions like social security systems influence care for the elderly. Such questions grow in significance as many regions transition from traditional to modern economies in the coming decades.
For example, Koda and Uruyos (2018) find that public transfers can crowd out private care for the elderly, and that the implications of modernization, including lower fertility and smaller families, tend to reduce familial support. These findings align with theories in Koda and Uruyos (2018) and Cigno (1993). However, Ko and Möring (2021), in their study of pensions in rural China, report that public pensions increase both well-being and intergenerational exchange. Similarly, Kodate and Timonen (2017) describe the complex interplay between public funding, paid healthcare workers, cultural norms, and family care.
The treatment of the elderly can influence and be influenced by consumption smoothing. Allocating resources for the future may enhance elderly care through cultural norms, asset accumulation, or both, depending on property rights and the nature of production.
Our model is constructed to shed light on these trade-offs, offering insight into how elderly treatment may evolve in the coming decades. By capturing the interaction between elderly treatment and economic attributes, the model generates predictions consistent with the complexity observed in the ethnographic record. In addition, it helps clarify vague concepts in the empirical literature such as ‘status,’ ‘prestige,’ or what it means for the elderly to be ‘treated well.’
@@ 4 Theoretical Framework
The model is a standard infinite-horizon overlapping-generations framework, where young, middle-aged, and elderly agents coexist at any given time. Decision-making occurs solely in middle age. Young agents derive no utility but may be inculcated with an appreciation for culture. In the base model, middle-aged individuals cultivate a preference in the young for giving gifts to the elderly. Upon reaching middle age, these individuals transfer resources to the elderly, deriving ‘warm glow utility’ as described by Andreoni (1989). This willingness to transfer depends on the degree of appreciation instilled in them during youth by the current middle-aged cohort. [^7] This approach resembles models in which cultural heritage and societal devotion are endogenously determined, such as Cozzi (1998). However, Cozzi’s model assumes self-fulfilling beliefs about culture without active cultivation.
@@
Middle-aged and elderly consumption at time $t$ are denoted $c_{mt}$ and $c_{et}$, with gifts to the elderly represented by $g_{mt}$. The lifetime utility of a middle-aged agent is given by the log-linear utility function:
$U_{t}=(1-\eta_{t})\ln c_{mt}+\eta_{t}\ln g_{mt}+\beta\ln c_{et+1},$ (1)
@@
where $\eta_{t}$ measures the relative importance of gift-giving in middle-aged utility.
Middle-aged agents receive income $y_{mt}$, allocated to consumption $c_{mt}$, savings $s_{mt}$, gifts $g_{mt}$, and inculcative investment in the young. Inculcation determines the future demand parameter $\eta_{t+1}$, reflecting the young’s appreciation for culture when they reach middle age. [^8] This contrasts with Olivera (2013) and Varvarigos (2021), where inculcation increases the likelihood of preference similarity across generations. These can be described as ‘Simpsonian’ models, inspired by Homer Simpson’s parenting philosophy: “Kids are the best… You can teach them to hate the things you hate.” (The Simpsons, season 11, episode 233, 1999). The cost of inculcation is $d_{t}\eta_{t+1}$, with $d_{t}=\delta y_{mt}$ ($\delta\in(0,1)$), scaling costs to income. [^9] This assumption ensures equilibria remain stationary, facilitating balanced-growth path analysis. Middle-aged consumption is thus:
$c_{mt}=y_{mt}-d_{t}\eta_{t+1}-s_{mt}-g_{mt}.$ (2)
Elderly consumption $c_{et}$ is derived from three sources: returns on previous savings $R_{t}s_{mt-1}$, earned income $y_{et}$, and received gifts $g_{et}$. Income $y_{et}$ may include compensation for productive or advisory roles, government transfers such as social security, or traditional sharing of output in hunter-gatherer societies. Elderly consumption is given by:
@@
This model of intergenerational gift-giving also captures cultural transmission. We will demonstrate that the utility function in (1) aligns with a model where cultural goods of endogenously determined value are exchanged across generations.
@@ 4.1 Simple model mechanics
We begin with a simplified version of the model in which there is no savings ($s_{mt}=0$ in Equation (3)). This simplification serves three purposes: it clarifies the intuition by focusing solely on the inculcation and gift decisions of current middle-aged agents; it empirically aligns with premodern economies where labor and natural resources, rather than capital accumulation, drive output—a characteristic of hunter-gatherer and subsistence agricultural societies; and it provides insight into the interplay between property rights, inculcation, and elderly treatment.
Substituting middle-aged consumption (2) and elderly consumption (3) into (1) and setting $s_{mt}=0$, we obtain:
@@
To simplify further, we assume $\eta_{t}\in{0,\eta}$ with $10<\eta<$ 1, so the current middle-aged generation makes a binary choice: whether to inculcate children to enjoy giving gifts to the elderly at intensity $\eta$ and cost $d_{t}$, or not. Simultaneously, they decide on a gift for the current elderly, based on the previous generation’s inculcation. These assumptions allow us to model the problem as a stationary, infinite-horizon game where each generation sequentially decides on inculcation, anticipating that future generations will adhere to the subgame-perfect Nash equilibrium strategy.
Middle-aged utility in an inculcation equilibrium is:
@@
Here, $g_{et+1}$ is the expected gift in the next period, given that current youth are inculcated. Stability of this equilibrium can be analyzed by considering a deviation where a middle-aged agent opts out of inculcation, leading to no future gifts. In this case, the agent’s lifetime utility becomes:
$U_{mt,0}=(1-\eta)\ln(y_{mt}-g^{\prime}{mt})+\eta\ln g^{\prime}{mt}+\beta\ln(y_{et+1})$ (6)
When inculcation and gift-giving are maintained, optimal gifts are derived by differentiating (5) with respect to $g_{mt}$, yielding:
@@
Thus, the middle-aged agent gives a gift proportional to disposable income, defined as income remaining after inculcation. Using this, equilibrium gifts received, $g^{et}$, can be determined. To account for demographic changes, we assume a constant population growth rate $n$, such that $n{t+1}=(1+n)n_{t}$. In equilibrium, this implies $g^{et}=(1+n)g^{*}{mt}$.
If the middle-aged agent abandons inculcation, the final gift given is $g^{\prime}{mt}=\eta y{mt}$, and no gift is received during old age in the subsequent period.